OpenStudy (anonymous):

Evaluate: ∫ x (1-x )^.5dx

OpenStudy (mimi_x3):

\[\large \int\limits x\sqrt{1-x}dx \] Let u = 1-x

OpenStudy (ash2326):

apmath did you get it?

OpenStudy (anonymous):

yeah that helped alot, thank you